A sample of 64 account balances from a credit company showed an average daily balance of $1,045. The standard deviation of the population is known to be $240. We are interested in determining if the mean of all account balances (i.e., population mean) is significantly different from $1,000.

Question:

Using the p-value approach at 95% confidence, test the above hypotheses.
Answer #1

Answer:

Given,

sample n = 64

population mean mu = 1000

sample mean = xbar = 1045

standard deviation = s = 240

alpha = 1 - 0.95

= 0.05

Here it is two tailed test

Ho : u = 1000

Ha : u != 1000

Now consider,

test statistic = (xbar - mu) / (s/sqrt(n))

substitute values

= (1045 - 1000) / (240/sqrt(64))

z = 1.5

P value = 2P(Z > z)

= 2P(Z > 1.5)

= 2*0.0668072 [since from z table]

P value = 0.1336

Here p > alpha(0.05), so we reject Ho.

So we conclude that there is enough evidence to claim that the population mean μ is different than 1000, at the 0.05 significance level.